OpenStudy (anonymous):

limit x->0 of x sins/cosx - 1

OpenStudy (anonymous):

I'd multiply by the conjugate first. \(\cos x +1 \)

OpenStudy (anonymous):

Quite a few ways to go about this.

OpenStudy (anonymous):

Well, if you use l'Hospital's rule it might be a lot easier.

OpenStudy (anonymous):

I tend to assume they haven't learned it.

OpenStudy (anonymous):

Is it 1-coax/x = 0 ? I just don't know how to apply it to this question

OpenStudy (anonymous):

Nope, first multiply top and bottom by \(\cos x+1\)

OpenStudy (anonymous):

Hmm, I was mistaken. I was thinking \(\sin x\approx x\) near 0, but that leads to indeterminate form again @wio

OpenStudy (anonymous):

Though you could use \(\sin x\approx x\) and have \[\frac{\sin^2x}{1-\cos x}=\frac{1-\cos^2x}{1-\cos x}\] Then factor, etc

OpenStudy (anonymous):

\[ \begin{split} \lim_{x\to 0} \frac{x\sin x}{\cos x - 1} &=\lim_{x\to 0} \frac{x\sin x(\cos x +1)}{\cos^2 x - 1}\\ &=\lim_{x\to 0} \frac{x\sin x(\cos x +1)}{-\sin^2 x}\\ &=\lim_{x\to 0} -\frac{x(\cos x +1)}{\sin x}\\ &=\lim_{x\to 0} -\frac{x}{\sin x}(\cos x+1)\\ &=\lim_{x\to 0} -\left(\frac{\sin x}{x}\right)^{-1} (\cos x+1) \end{split} \]

OpenStudy (anonymous):

Wtf fractions aren't working anymore for me.

OpenStudy (anonymous):

\[ \frac{1}{2} \]

OpenStudy (anonymous):

\[\frac{ a }{ s }\]